the length of a rectangle is 1 centimetre more than its breadth if the sum of length and breadth is 2 centimetre more than its diagonal then find the length of a rectangle
Answer by jorel1380(3719)

Let n be the breadth of the rectangle. Then it's length is n+1. So:
n+n+1-2=√((n²)+(n+1)²)
2n-1=√((n²)+(n+1)²)
4n²-4n+1=n²+n²+2n+1
2n²-6n=0
2n²=6n
n²=3n
n=3
n+1=4
The rectangle is 4 cms. in length. ☺☺☺☺